Problem Set 3 - Chp. 4 In Exercise 1, find the indicated binomial probabilities. 1f convenient, use technology or Table 2 in Appendix B of the e-text. 1. Sixty-two percent of U.S. adults get news on social media sites. You randomly select five U.S adults. Find the probability that the number of U.S. adults who get news on social media sites is (a) exactly two, (b) at least two, and (c) more than two. (Source: Pew Research Center) In Exercise 2, find the mean, variance, and standard deviation of the binomial distribution for the given random variable. Interpret the results 2. About 13% of U.S. drivers are uninsured. You randomly select eight US. drivers and ask them whether they are uninsured. The random variable represents the number of U.S. drivers who are uninsured. (Source: Insurance Research Council) In Exercises 3-5, find the indicated probabilities using the geometric distribution, the Poisson distribution, or the binomial distribution. Then determine whether the events are unusual. If convenient, use a table or technology to find the probabilities 3. Thirty-six percent of Americans think there is still a need for the practice of changing their clocks for Daylight Savings Time. You randomly select seven Americans. Find the probability that the number of Americans who say there is still a need for changing their clocks for Daylight Savings Time is (a) exactly four, (b) less than two, and (c) at least six. (Source: Rasmussen Reports) 4. In a recent season, hockey player Evgeni Malkin scored 27 goals in 57 games he played Assume that his goal production stayed at that level for the next season. Find the probability that he would get his first goal (a) in the first game of the season, (b) in the second game of the season, (c) within the first three games of the season, and (d) not within the first three games of the season. (Source: National Hockey League) 5. During a 10-year period, sharks killed an average of 6.1 people each year worldwide. Find the probability that the number of people killed by sharks next year is (a) exactly three, (b) more than six, and (c) at most five. (Source: International Shark Attack File)
Answer #1

1)

X follow binomial distribution with p = 0.62 and n = 5
a)
P(X = 2) = 0.2109
b)
P(X >= 2) = 0.9274
c)
P(X > 2) = 0.7165

2)
mean = np = 8 * 0.13 = 1.04
sd =sqrt(npq) = sqrt(8 *0.13 * (1 - 0.13)) =   0.95120975604

Please post rest questions again